OpenStudy (anonymous):

Help me with multiplying monomials? (3x^3y^4)^2 I got 3x^6y^8 Is that right?

OpenStudy (anonymous):

9 instead of 3

OpenStudy (anonymous):

do you mean (3x(multiplied) by 3y (multiplied) by 4) all multiplied by 2? * is the usually sign for multiplication.

OpenStudy (anonymous):

so do I multiply everything by the 2?

OpenStudy (anonymous):

 you multiply the 2 with each #. 2*3x= 6x, 2*3y=6y, 2*4=8